1. In the context of the CAPM, the relevant risk is: a. Unique Risk b. Market risk c. Standard deviation of returns. d. Variance of returns e. Derivatives risk. 2. According the CAPM a well-diversified portfolios rate of return is a function of: a. Systematic risk b. Unsystematic risk c. Unique risk d. Reinvestment risk e. Credit risk f. Derivatives risk 3. The RF rate and the expected market rate of return are 6 and 12% respectively. According the CAPM, the expected rate of return on Stock X with a beta of 1.2 is: a, b. c. d. e. f. 6% 14.4% 12% 13.2% 18% 12.9%
Answer #1

1.In the context of CAPM, the relevant risk is none of the above.

Beta is the relevant risk in the context of CAPM.

2.According to CAPM, a well-diversified portfolio’s rate of return is a function of systematic risk.

In a well diversified portfolio, the only return that affects CAPM is beta.

3.Given:

Risk free rate= 6%

Marker rate of return= 12%

Beta= 1.2

The expected rate of return is calculated using the formula below:

Ke=Rf+\beta[E(Rm)-Rf]

Where:

Rf=risk-free rate of return

Rm=expected rate of return on the market.

\beta= stock’s beta

Ke= 6+1.2(12-6)

     = 6+ 7.2= 13.2%.

Therefore, the expected rate of return is 13.2%.

I hope that was useful :)